You will often see athletes being treated with disposable cold packs if they injure themselves in an event. These cold packs contain water and ammonium nitrate. Ammonium nitrate has an enthalpy of solution of 25.7 kJ/Mol. Determine the temperature change of solution if 45.0g of ammonium nitrate is dissolved in 150ml of water, and thus the temperature of the cold pack if it's initial temperature is 20.0oC. Assume that the specific heat of the solution is the same as water 4.186.

1
Expert's answer
2021-09-10T00:40:41-0400

∆H = m x s x ∆T

25700 = 195 × 4.186 (X -293)

X - 293 = 25700/816.27

X - 293 = 31.48

X = 293 + 31.48 = 324.48

=51.48°C